MICROWAVE SPECTROSCOPY OF TWO ISOELECTRONIC RADICALS: 2,4,6-HEPTATRIYNYL, H_{2}C-C≡C-C≡C-C≡C-H, AND 2,4-HEXADIYNYLNITRILE, H_{2}C-C≡C-C≡C-C≡N

WEI CHEN, Harvard-Smithsonian Center for Astrophysics, Cambridge, MA 02138; and Division of Engineering and Applied Sciences, Harvard University, Cambridge, MA 02138; and Department of Chemistry, Wesleyan University, Middletown, CT 06459; \underline{STEWART E. NOVICK}, Department of Chemistry, Wesleyan University, Middletown, CT 06459; M. C. McCARTHY, P. THADDEUS, Harvard-Smithsonian Center for Astrophysics, Cambridge, MA 02138; and Division of Engineering and Applied Sciences, Harvard University, Cambridge, MA 02138.

The isoelectronic pair of radicals 2,4,6-heptatriynyl, H_{2}C-C≡C-C≡C-C≡C-H, and 2,4-hexadiynylnitrile, H_{2}C-C≡C-C≡C-C≡N, were produced in a pulsed discharge supersonic jet and detected by Fourier transform microwave spectroscopy. Seven rotational transitions of the heptatriynyl radical from N′ = 4 to 10 and six rotational transitions of the hexadiynylnitrile radical from N′ = 4 to 9 in the K_{a} = 0 ladder were observed between 6 and 17 GHz. For each species, over 50 hyperfine components were measured and assigned with an uncertainty of 5 kHz. Precise rotational, centrifugal distortion, spinrotation, and various hyperfine coupling constants were determined for both H_{2}C_{7}H and H_{2}C_{6}N.